OpenStudy (anonymous):

Medal award! What are the positive and negative measures of the closest coterminal angles to 1.1 radians?

zepdrix (zepdrix):

Hmm I don't understand the question :d Do they want the degree measure? Or they want the closest "special angle" maybe?

OpenStudy (anonymous):

I think they want radians

zepdrix (zepdrix):

Hmm ok :) Then for co-terminal we want to spin a full rotation around the circle and land in the same spot. So one co-terminal angle would be: \(\large\rm 1.1+2\pi\)

zepdrix (zepdrix):

How bout if we want to spin backwards? Any ideas? :)

OpenStudy (anonymous):

the thing is I believe they want us to convert the 1.1 into a fraction

OpenStudy (anonymous):

and of course you will go clockwise for the negative :)

zepdrix (zepdrix):

Hmm there is no reason for us to write as a fraction if we're staying in radians. Unless you want to relate it to one of your special angles. 1.1 radians is approximately 31.5 degrees which is approximately pi/6. But I don't think that's what they're asking for. Hmmm....

OpenStudy (anonymous):

I put in 11pi/10-10pi/5 which is -9pi/10 for the negative angle. and for the positive 11pi/10+10pi/5 which is 31pi/10 but that is incorrect.

zepdrix (zepdrix):

Well keep in mind that:\[\large\rm 1.1=\frac{11}{10}\]\[\large\rm 1.1\ne\frac{11\pi}{10}\]

zepdrix (zepdrix):

You can get a common denominator if you like,\[\large\rm \frac{11}{10}+\frac{20\pi}{10}=\frac{11+20\pi}{10}\]But it looks a lot nicer if you just leave it as \(\large\rm 1.1+2\pi\)

zepdrix (zepdrix):

Just don't combine the numerators:\[\large\rm 11+20\pi\ne31\pi\]:)

OpenStudy (anonymous):

so for the positive angle it would just be 1.1+2pi?

zepdrix (zepdrix):

ya i think that's all they want us to input. Do you get multiple guesses? XD

OpenStudy (anonymous):

I get 3 attempts and then I get a new problem

OpenStudy (anonymous):

and for the negative it was be 1.1-2pi?

zepdrix (zepdrix):

yes

OpenStudy (anonymous):

that's correct :) thanks so much

zepdrix (zepdrix):

yay team \c:/

OpenStudy (anonymous):

:D